OpenStudy (commonlizer):

Which of these is a simplified form of the equation 7p + 4 = − p + 9 + 2p + 3p? A 13p = 13 B 3p = 5 C 7 = 4 D 11 = 15

OpenStudy (commonlizer):

my answer is C

OpenStudy (commonlizer):

am I write

Mehek (mehek14):

Wrong

OpenStudy (commonlizer):

really I stink in math plz hlp me

Mehek (mehek14):

7p + 4 = − p + 9 + 2p + 3p 7p + 4 = 9 + 4p now move 4 to the other side and 4p to the left

OpenStudy (commonlizer):

okay I got it

Mehek (mehek14):

So what would the answer be?

OpenStudy (commonlizer):

11=13 right

Mehek (mehek14):

nope 9 - 4 = 5 7p - 4p = 3p 3p = 5 I asked to move 4 to the other side and 4p to the left

OpenStudy (alex136):

\[7p+4=−p+9+2p+3p \] Step 1: Simplify both sides of the equation. \[7p+4=−p+9+2p+3p \] Simplify: \[7p+4=(−p+2p+3p)+(9) \](Combine Like Terms) \[7p+4=4p+9 \] Step 2: Subtract 4p from both sides. \[7p+4−4p=4p+9−4p \] \[3p+4=9 \] Step 3: Subtract 4 from both sides. \[3p+4−4=9−4 \] \[3p=5 \]

OpenStudy (commonlizer):

oh srry wait let write this down

OpenStudy (commonlizer):

okay thank you both I have taken this down

OpenStudy (alex136):

Your welcome